OpenStudy (anonymous):

Initially sliding with a speed of 2.2m/s, a 1.8kg block collides with a spring and compresses it 0.31m before coming to rest. What is the force constant of the spring?

OpenStudy (anonymous):

Assuming energy is conserved Ek = Ep(sring) 1/2*m*v^2 = 1/2*k*x^2 1/2(1.8)*((2.2)^2) = 1/2*k*((0.31)^2) 4.356 = 0.04805*k k = 90.66 (2dp) N/m

OpenStudy (anonymous):

the kinetic energy possesed by the body is converted to spring energy as per the above equation